The Warren Watch Company sells watches for $30, fixed costs are $180,000, and variable costs are $13 per watch.

  1. What is the firm's gain or loss at sales of 5,000 watches? Loss, if any, should be indicated by a minus sign. Round your answer to the nearest cent.
    $ _____
  2. What is the firm's gain or loss at sales of 15,000 watches? Loss, if any, should be indicated by a minus sign. Round your answer to the nearest cent.
    ____ $  
  3. What is the break-even point (unit sales)? Round your answer to the nearest whole number.
    ____ units
  4. What would happen to the break-even point if the selling price was raised to $33?   The result is that the break-even point increases. The result is that the break-even point decreases. The result is that the break-even point remains unchanged.
  5. What would happen to the break-even point if the selling price was raised to $33 but variable costs rose to $24 a unit? Round your answer to the nearest whole number.
    The result is that the break-even point increases. The result is that the break-even point decreases. The result is that the break-even point remains unchanged.
Answer #1

1.
=5000*(30-13)-180000
=-95000

2.
=15000*(30-13)-180000
=75000

3.
=180000/(30-13)
=10588.23529

4.
The result is that the break-even point decreases.
=180000/(33-13)
=9000

5.
The result is that the break-even point increases.
=180000/(33-24)
=20000